OpenStudy (anonymous):

Set two planes whose intersection is the parametric line x=2+t; y=3t; z=1-t (t belongs to R)

OpenStudy (dumbcow):

The line has a directional vector of <1,3,-1> that goes through point (2,0,1) The 2 planes will be in this form: ax +by +z = c dx +ey +z = f the normal vectors are <a,b,1> and <d,e,1> respectively by taking the cross product of the normal vectors you can obtain the vector <1,3,-1> also plugging in the point (2,0,1) into both plane equations for x,y,z this yields the following 5 equations: b-e = 1 a-d = -3 ae -bd = -1 2a +1 = c 2d +1 = f since there are 6 variables, 1 is a free variable. Let f = 0 --> d = -1/2 --> a = -7/2 --> c = -6 this leaves b-e = 1 -7/2e +1/2b = -1 --> e = 1/2 --> b = 3/2 So 2 possible planes (written in standard form) that intersect with given line are -7x +3y +2z = -12 -x + y +2z = 0